7)f(x)= 10 -x^3-3cosx, estimate therootusing secant method up to 3 significant
Answer #1

given fra)= 10-13-30054 7 0 2 3 7.3751 3.24841-14-03 Root Is lies between 2013 1st Iteration 2-188 let ao=2:21=3 fero)=f(2)=and Iteration 4 = 2-688 : 213=2255 f(n) = f(2-18€) = 1.2615; $(33) = $12-255)= 0.4295 A4 = 22 f(na)– 13 frag) 2-168(0.4295)-2